Ticket Number: HEA-NOC/20080811-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20080811 09:52 UTC+1 Problem Start: 20080811 09:27 UTC+1
Ticket Update: 20080811 12:31 UTC+1 Problem End: 20080811 09:29 UTC+1
NCAD Loss of connectivity
Brief loss of connectivity
20080811 09:44 COC Contacted NCAD, awaiting call back
20080811 10:28 COC Circuit has gone down again. Contacted BT and NCAD.
NCAD have confirmed they are not carrying out any maintenance. Opened
ticket with BT
20080811 11:05 COC Circuit has gone down again. Contacting BT again
20080811 11:18 COC Problem is between Fenian street and NCAD. NCAD are
double checking their cabling. BT has placed a soft loop at Fenian street.
20080811 12:19 COC NCAD have determined that these outages were caused by
contractors onsite tripping the power supply to the comms room housing the
cpe device. Monitoring for 24 hours
20080811 12:29 COC NCAD will be moving this equipment onto another circuit
at 08:30 tomorrow which will cause another outage.
This ticket can be monitored at http://www.hea.net/tickets/20080811-1
HEAnet Limited. Registered in Ireland, No. 275301.